A solution is prepared by dissolving 10.00 g sucrose (MM of sucrose = 343 g) in 50g of water. The volume of the resulting solution is 52 mL. Calculate the following

A. molarity

B. molality

C. mole fraction


1
Expert's answer
2022-03-08T02:00:04-0500

nsucrose = 10/343 = 0.0291 mol

nwater = 50/18 = 2.78 mol


Molarity or molar concentration is the number of moles of solute per liter of solution:


csucrose = nsolute/Vsolution =1000*0.0291/52 = 0.56 mol/l


Molality is the number of moles of solute per kilogram of solvent:


bsucrose = nsolute/msolution = 1000*0.0291/50 = 0.582 mol/kg


Mole fraction is the number of molecules (or moles) of one component divided by total the number of molecules (or moles) in the mixture:


xsucrose = 0.0291/(0.0291+2.78) = 0.0104